OpenStudy (anonymous):

solve for x: 2/x-2+7/x^2-4=5/x

OpenStudy (anonymous):

Can you use the equation editor so your equations are a little less ambiguous?

OpenStudy (mimi_x3):

is it: \[\frac{2}{x-2}+ \frac{7}{x^{2}-4} = \frac{5}{x} \]

OpenStudy (anonymous):

use brackets plz

OpenStudy (anonymous):

x=5 or x=-4/3,